2. a. Explain in words how and why the multiplier aG and the interest sensitivity of aggregate
demand affect the slope of the IS curve.
b. Explain why the slope of the IS curve is a factor in determining the working of monetary
policy.
1
Expert's answer
2021-08-09T06:49:31-0400

The slope of the IS curve. The slope of the IS curve is 1 / (kAb) or MLR / b, where MLR is the maximum seizure rate (recall that MLR = 1 - mpc (1 - t) + mpm = mps (1 - t) + t + mpm, i.e. That is, the marginal rate of withdrawals is the reciprocal of the multiplier of expenditures, MLR = 1 / kA). Thus, the slope of the IS curve is determined by: 1) the sensitivity of autonomous expenses to the interest rate (b),


2) the value of the multiplier (kA), which depends on the marginal propensity to consume (mpc), the tax rate (t) and the marginal propensity to import (mpm).


The slope of the IS curve decreases (it turns clockwise and becomes flatter). The IS curve will be flatter:


The sensitivity of autonomous expenditures to the interest rate (b) is high, which


means that even a small change in the interest rate leads to a significant change in autonomous costs and, therefore, income;


· The expense multiplier (kA) is large, and the marginal withdrawal rate (MLR) is small, which is possible if: a) the marginal propensity to consume is high; b) the marginal tax rate is low; c) the marginal propensity to import is small. If the multiplier is large, it means that even a minor change in autonomous spending will lead to a large multiplicative change in income. (Note that the magnitude of the multiplier determines both the slope and the magnitude of the shift of the IS curve).


Thus, increasing b and mpc and decreasing t and mpm decrease the slope of IS.


The slope of the IS curve increases (it turns counterclockwise and becomes steeper) as b and/or kA decreases


The IS curve, however, does not determine either a specific value of the level of income Y, or a single value of the equilibrium interest rate R, it only reflects all possible combinations of Y and R, in which the market for goods and services is in equilibrium. Therefore, in order to determine their values, one more equation with the same variables is needed. To do this, you should turn to the money market.